OpenStudy (anonymous):

solve using substitution method 10x + y = –1 5x + 2y = 16

OpenStudy (saifoo.khan):

x = -6/5 y = 11

OpenStudy (anonymous):

solve for y on the first equation \[y = -10x - 1\] Substitute y into the second equation \[5x + 2(-10x - 1) = 16\] Simplify and solve for x\[5x - 20x - 2 = 16\] \[-15x = 18\] \[x = - \frac{6}{5}\] Plug in - 6/5 into any one of the original equations and you get \[10(-\frac{6}{5}) + y = -1\]\[-12 + y = -1\] y = 11 So the solution is \[(-\frac{6}{5},11)\]